(1) set the minimum support criterion at 0.6, identify the frequent 2-itemsets;

Transaction Item Purchased
1 A B C
2 A C D
3 B C D
4 A D E
5 B C E
1 item set Support
2 itemset Support
Frequent 2 Itemsets

calculate lift ratio for below rules and interpret the results

Rule Support of Antecedent {a} Support of Consequent {c} Support of {a & c} The confidence of the Rule Lift Ratio of the Rule Interpretation on Lift Ratio
A-->D
C-->A
A-->C
B&C-->D
Answer #1

According to Association , the minimum support = number of transaction * support criterion = 5 * 60 /100 = 3

Item set = A B C D E

Support count (\sigma) = frequency of an item set

  • \sigma(B,C) = 3 (* B and C occur thrice in the 5 transactions )
  • \sigma(A,C)= 2
  • \sigma(C,D)= 2

Hence, the three minimum support and frequent data sets are B,C ; A,C ; C,D

Calculating the lift ratio

Rule Support of Antecedent {a} Support of Consequent {c} Support of {a & c} The confidence of the Rule Lift Ratio of the Rule
A-->D 3/5 = 0.6 3/5 = 0.6 2/5 = 0.4 \sigma(A,D)/ \sigma(A) = 0.67 Confidence(A,D)/ Expected confidence(D) = 40% /60%= 0.67
C-->A 4/5= 0.8 3/5 = 0.6 2/5 = 0.4 0.5 0.4/0.6 = 0.67
A-->C 3/5 = 0.6 4/5=0.8 2/5= 0.4 0.67 0.4/ 0.8 = 0.5
B&C-->D 1/5=0.2 3/5 = 0.6 1/5= 0.2 \sigma(B,C,D)/\sigma(B,C)= 1 0.2/0.6= 0.33

( *The below value means that the number of transaction occurance of A->D and the total transactions  of A.

Rule

Support of Antecedent {a}
A-->D 2/4 = 0.5

)

Based on the Lift ratio,

  • The probability of person taking A will take D & the person taking C will take A is 67 %
  • The probability of person taking A will take C is 50%
  • The probability of person taking B&C will take D is 33%

The life ratio implies the relationship between the antecedent and consequent. The probability will increase with the increase in the lift ratio.